xmoseje:

Students should be able to find sine, cosine, tangent in a right triangle given 3 side lengths. They should also remember the side ratios of the special triangles and find s/c/t. They should also be able to identify opposite and adjacent sides. find sin (a) 7/25 24/7 24/25

xmoseje:

@bonez

xmoseje:

@Elsa213

Shadow:

@xmoseje We need a picture of the triangle in order to do this.

xmoseje:

1 attachment
xmoseje:

i forgot

Shadow:

But the way to remember how to identify sin, cos, and tan is: SOH CAH TOA \[\sin = \frac{ opposite }{ hypotenuse }\] \[\cos= \frac{ adjacent }{ hypotenuse }\] \[\tan = \frac{ opposite }{ adjacent }\]

xmoseje:

thnx

Shadow:

sin(a) so start at angle A then look at the opposite angle and put it over the hypotenuse.